# Uncertainty-Based Extensible Codebook for Discrete Federated Learning in Heterogeneous Data Silos

### v1

@@ 4.1 Extensible Codebook
Number of Codewords. We investigate the impact of varying the number of initialized codewords within our extensible codebook, aiming to strike a balance between achieving competitive accuracy and optimizing the runtime efficiency of the K-means initialization. Our findings, illustrated in Figure 4 for the GTSRB dataset, reveal that starting with 32 or 64 codewords offers comparable accuracy and uncertainty metrics to larger codebooks, while significantly enhancing the efficiency of the K-means initialization. This efficiency highlights the efficacy of our proposed approach.
Figure 4: For our extensible codebook, 32 or 64 initialized codewords achieve competitive performance.
@@
Extensible Codebook v.s. Static Large Codebook. To validate our extensible codebook’s superiority over starting with a large codebook, we ensured both methods ended with the same number of codewords through experiments. For the CIFAR100 dataset, the extensible codebook was initially set to 128 codewords and expanded twice, while the static codebook was fixed at 512 codewords. Results showcased in Table 1 demonstrate the difficulties associated with a larger initial codebook in codeword selection for image features. Conversely, gradually expanding the codebook significantly improved codeword differentiation, yielding better outcomes, such as enhanced accuracy (0.375 for $\mathcal{D}_{1}$) and reduced uncertainty (0.78 vs. 1.66 for the static approach). In addition, perplexity results reveal increased utilization of our extensible codebook, offering clear evidence of our design’s superiority.
Codebook Data $\mathbf{\mathcal{L}_{code}}$ mP mE mA
w/o init $\mathcal{D}_{1}$ 0.6202 6.26 0.125 0.888
$\mathcal{D}_{2}$ 0.6063 5.99 0.296 0.554
$\mathcal{D}_{3}$ 0.6096 5.35 0.267 0.622
w/ init $\mathcal{D}_{1}$ 0.0862 59.64 0.0935 0.945
$\mathcal{D}_{2}$ 0.0785 57.58 0.1604 0.906
$\mathcal{D}_{3}$ 0.0779 99.38 0.1509 0.929
Table 3: Comparison with different methods on heterogeneous data. mA denotes the mean of Accuracy for all data silos from the same distribution, and mE denotes the mean of Entropy. DisTrans lacks a Dropout layer, rendering it incapable of evaluating uncertainty.
Methods Data MNIST FMNIST GTSRB CIFAR10 CIFAR100
mA mE mA mE mA mE mA mE mA mE
Fedavg(McMahan et al., 2017) $\mathcal{D}_{1}$ 0.874 0.212 0.801 0.246 0.670 0.623 0.676 0.172 0.110 1.74
All 0.780 0.273 0.803 0.273 0.660 0.636 0.617 0.177 0.088 1.91
DisTrans(Yuan et al., 2022) $\mathcal{D}_{1}$ 0.856 - 0.721 - 0.898 - 0.721 - 0.289 -
$\mathcal{D}_{2}$ 0.799 - 0.705 - 0.900 - 0.719 - 0.261 -
$\mathcal{D}_{3}$ 0.789 - 0.694 - 0.897 - 0.659 - 0.251 -
All 0.815 - 0.707 - 0.898 - 0.699 - 0.267 -
All 0.920 0.149 0.850 0.167 0.942 0.0239 0.720 0.0222 0.326 0.655
4.2 UEFL for Heterogeneous Datasets
To underscore the dual enhancement of accuracy and uncertainty reduction by our UEFL, we conducted comparative experiments on five renowned datasets against leading algorithms, specifically the baseline Federated Averaging (FedAvg)(McMahan et al., 2017) and DisTrans(Yuan et al., 2022). We structured the evaluation around 9 data silos, derived from three distinct data distributions as delineated in Section 4, with every three data silos originating from a single distribution. For accuracy comparison, DisTrans generally exhibits better performance than FedAvg, making it our primary point of comparison. However, since DisTrans works differently, we show the results for the average results not only for data distribution $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, and $\mathcal{D}_{3}$ but also for all data as shown in Table 3. Regarding uncertainty comparison, because DisTrans lacks Dropout layers, precluding uncertainty evaluation, we exclusively compare uncertainty metrics with FedAvg.
Accuracy. The results in Table 3 provide a comprehensive comparison, illustrating that our UEFL surpasses all other state-of-the-art (SOTA) methods in both accuracy and uncertainty reduction. Specifically, UEFL improves accuracy over FedAvg by 17.94% and DisTrans by 12.88% for the $\mathcal{D}_{3}$ distribution of the MNIST dataset. Figure 5 details performance across individual data silos, highlighting our UEFL’s effectiveness in elevating the last three silos to parity with their counterparts. Overall, our UEFL achieves accuracy improvements ranging from 3% to 22.1% over DisTrans.
Uncertainty. The data in Table 3 highlight that our UEFL also excels in reducing uncertainty. For the MNIST dataset’s $\mathcal{D}_{3}$ distribution, our approach reduces uncertainty by 45.42%. The visualization in Figure 5 effectively demonstrates how our UEFL identifies the last three data silos from an unknown distribution, as their uncertainty is notably higher than that of others. Furthermore, the figure demonstrates how our UEFL reduces uncertainty by assigning new codewords, with beneficial effects extending to other distributions. Our UEFL improves uncertainty compared to FedAvg, achieving reductions by 38.83%-96.24%.
Figure 5: Detailed comparison for all data silos. Experiments are on MNIST. $\mathcal{D}_{3}$ presents much lower accuracy and higher uncertainty compared to $\mathcal{D}_{1}$, $\mathcal{D}_{2}$ for FedAvg. And the perplexity results show that our UEFL assigns new codewords to $\mathcal{D}_{3}$ to improve the performance.
Codewords Perplexity. Figure 5 also presents a perplexity comparison between our UEFL and FedAvg, illustrating enhanced codebook utilization after assigning new codewords to $\mathcal{D}_{3}$. This adjustment not only benefits $\mathcal{D}_{3}$ but also improves the codebook utilization for $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$.
4.3 Ablation Study
The Influence of Discretization. The codebook serves to distinctly differentiate data distributions, addressing heterogeneity. However, mapping all feature vectors to specific codewords can restrict the diversity of image representations, which may affect performance, particularly with a smaller initial codebook size. As shown in Table 4 for the GTSRB dataset, although discretization of DisFL (i.e. Discrete FL) leads to a minor performance decrease, our UEFL still delivers enhanced performance overall.
Table 4: The performance of DisFL drops because of the discretization of features to a small codebook. However, with the extensible capacity of our codebook, the final performance will be better.
Metric Data FL DisFL UEFL
mA $\mathcal{D}_{1}$ 0.897 0.712 0.95
$\mathcal{D}_{2}$ 0.851 0.673 0.964
$\mathcal{D}_{3}$ 0.803 0.602 0.911
Unbalanced Datasets. To showcase our method’s efficacy in managing unbalanced scenarios, we constructed an experimental setup with three data silos from $\mathcal{D}_{1}$ and one each from $\mathcal{D}_{2}$ and $\mathcal{D}_{3}$, totaling five silos. Table 5 demonstrates that our UEFL significantly improves both accuracy and uncertainty in such unbalanced data configurations.
Table 5: Our UEFL improves the performance of unbalanced data.
Data Silo Fedavg UEFL
Acc Entropy Acc Entropy
$\mathcal{D}_{1}$ $s_{1_{a}}$ 0.964 0.0312 0.952 0.0291
$s_{1_{b}}$ 0.936 0.0252 0.974 0.0170
$s_{1_{c}}$ 0.964 0.0499 0.944 0.0308
$\mathcal{D}_{2}$ $s_{2_{a}}$ 0.796 0.1477 0.836 0.1261
$\mathcal{D}_{3}$ $s_{3_{a}}$ 0.508 0.2560 0.828 0.1048
Segmented Codewords. For more complex datasets, requiring a broader representation of image features but with minimal initialization time, we employ codeword segmentation to enhance selection capacity efficiently. In Table 6, we explore the impact of segmentation factors of 1, 2, and 4, starting with 16 codewords for GTSRB and 32 for CIFAR100. Our findings indicate that, particularly for CIFAR100, splitting vectors into 4 segments with only 32 initialized codewords achieves impressive performance. Similarly, for GTSRB, segmentation into 2 parts is adequate for effective image feature representation.
Table 6: Comparison of different segments. The number of initialized codewords for GTSRB is 16, and for CIFAR100 is 32. ”-” denotes value close to 0.
Segment Data GTSRB CIFAR100
mA mE mA mE
s=1 $\mathcal{D}_{1}$ 0.188 2.81 2.86 2.75
$\mathcal{D}_{2}$ 0.183 2.61 0.953 2.72
$\mathcal{D}_{3}$ 0.206 2.63 0.119 2.53
s=2 $\mathcal{D}_{1}$ 0.918 - 0.362 1.97
$\mathcal{D}_{2}$ 0.903 - 0.335 2.04
$\mathcal{D}_{3}$ 0.863 - 0.274 2.18
s=4 $\mathcal{D}_{1}$ 0.923 - 0.401 0.821
$\mathcal{D}_{2}$ 0.911 - 0.351 0.853
$\mathcal{D}_{3}$ 0.862 - 0.305 0.881
@@ 4.3 Ablation Study
5 Conclusion
In this work, we address the challenge of data heterogeneity within federated learning by introducing an innovative solution: an extensible codebook designed to map distinct data distributions using varied codeword pools. Our proposed framework, Uncertainty-Based Extensible-Codebook Federated Learning (UEFL), leverages this extensible codebook through an iterative process that adeptly identifies data from unknown distributions via uncertainty evaluation and enriches the codebook with newly initialized codewords tailored to these distributions. The iterative nature of UEFL, coupled with efficient codeword initialization using K-means, ensures codewords are closely matched with the actual data distribution, thereby expediting model convergence. This approach allows UEFL to rapidly adjust to new and unseen data distributions, enhancing adaptability. Our comprehensive evaluation across five prominent datasets showcases UEFL’s effectiveness, yielding accuracy enhancements ranging from 3% to 22.1% and significant reductions in uncertainty between 38.83% and 96.24%.
Impact Statement
Appendix A Appendix
For the experiments on the CIFAR10 dataset, in Figure 7, at round 40, after we assign new codewords, rapid performance gains are evident. Remarkably, the training process demonstrates swift convergence, typically within just five rounds. For illustrative clarity and to underscore the differential impact, we extend the training to 20 rounds in subsequent iterations, showcasing the accelerated and effective adaptation of our approach. In addition, after 60 rounds, even if we keep adding new codewords, the increased perplexity denotes a higher utilization of the codebook. However, there is no significant improvement in accuracy or uncertainty.
Figure 7: Learning curves of CIFAR10. (left) Accuracy results. (mid) Uncertainty results. (right) codeword perplexity results.
For the experiments on the FMNIST and GTSRB datasets, as depicted in Figure 8 and Figure 9, we introduce new codewords to the codebook only once. Notably, there is a clear ”performance jump” evident in all six figures, showcasing the rapid adaptation of our UEFL to new data distributions. And Figure 10 shows a large jump for MNIST dataset with our UEFL.
Figure 8: Learning curves of FMNIST. (left) Accuracy results. (mid) Uncertainty results. (right) codeword loss results.
Figure 9: Learning curves of GTSRB. (left) Accuracy results. (mid) Uncertainty results. (right) codeword loss results.
Figure 10: Learning curves of MNIST.
In Figure 11, we explore the impact of different thresholds in our UEFL. When the threshold $\gamma$ is set to 0.2, 0.3, or 0.4, it proves to be too high, and all data silos easily meet the criterion. Consequently, our UEFL doesn’t add any codewords to the codebook, resulting in no performance improvement. It’s only when we lower the threshold $\gamma$ to 0.1 that our UEFL effectively identifies the heterogeneous data distribution and allocates new codewords to enhance performance. This aligns with the findings presented in Table 7.
Figure 11: Learning curves of CIFAR100. (Left Top) $\gamma=0.1$. (Right Top) $\gamma=0.2$. (Left Bottom) $\gamma=0.3$. (Right Bottom) $\gamma=0.4$.
$\gamma$ Data $\mathcal{L}_{code}$ ppl unc acc
0.5 $\mathcal{D}_{1}$ 3.57 14.84 0.814 0.575
$\mathcal{D}_{2}$ 3.55 15.37 0.780 0.657
$\mathcal{D}_{3}$ 3.25 15.01 0.821 0.605
0.4 $\mathcal{D}_{1}$ 3.47 15.49 0.681 0.647
$\mathcal{D}_{2}$ 3.60 15.56 0.672 0.674
$\mathcal{D}_{3}$ 3.41 14.99 0.732 0.582
0.3 $\mathcal{D}_{1}$ 3.52 17.01 0.622 0.652
$\mathcal{D}_{2}$ 3.67 16.87 0.614 0.693
$\mathcal{D}_{3}$ 3.44 16.47 0.645 0.639
0.2 $\mathcal{D}_{1}$ 1.53 24.24 0.0843 0.936
$\mathcal{D}_{2}$ 1.62 23.99 0.0896 0.941
$\mathcal{D}_{3}$ 1.64 25.84 0.167 0.881
0.1 $\mathcal{D}_{1}$ 0.387 25.85 0.0196 0.95
$\mathcal{D}_{2}$ 0.346 25.94 0.0206 0.964
$\mathcal{D}_{3}$ 0.458 26.96 0.0314 0.911
Table 7: Uncertainty Evaluator threshold.Results on GTSRB dataset with 64 initialized codewords with segment 1.
As demonstrated in Table 8, the GTSRB dataset highlights that our EFL requires just 64 initialized codewords to achieve competitive performance.
#Codes Data $\mathbf{\mathcal{L}_{code}}\downarrow$ mPnormal-$\uparrow$ mEnormal-$\downarrow$ mAnormal-$\uparrow$
8 $\mathcal{D}_{1}$ 3.61 4.78 2.02 0.257
$\mathcal{D}_{2}$ 3.68 4.86 2.01 0.265
$\mathcal{D}_{3}$ 3.38 4.81 2.03 0.249
16 $\mathcal{D}_{1}$ 3.41 10.46 1.36 0.515
$\mathcal{D}_{2}$ 3.55 10.54 1.37 0.506
$\mathcal{D}_{3}$ 3.23 10.61 1.39 0.486
32 $\mathcal{D}_{1}$ 0.127 25.91 0.0412 0.956
$\mathcal{D}_{2}$ 0.0885 25.42 0.0313 0.966
$\mathcal{D}_{3}$ 0.178 26.37 0.117 0.911
64 $\mathcal{D}_{1}$ 0.0975 26.79 0.0086 0.965
$\mathcal{D}_{2}$ 0.0853 26.32 0.0084 0.974
$\mathcal{D}_{3}$ 0.1907 27.23 0.0166 0.926
128 $\mathcal{D}_{1}$ 0.0512 38.73 - 0.954
$\mathcal{D}_{2}$ 0.0453 34.16 - 0.968
$\mathcal{D}_{3}$ 0.0726 43.42 - 0.917
256 $\mathcal{D}_{1}$ 0.0543 41.20 0.0043 0.962
$\mathcal{D}_{2}$ 0.0301 38.96 0.0054 0.959
$\mathcal{D}_{3}$ 0.0577 50.57 0.0103 0.904

### v4

@@ Appendix I Our UEFL for Label Heterogeneity
Number of Codewords. We investigate the impact of varying the number of initialized codewords within our extensible codebook in Table 14, aiming to strike a balance between achieving competitive accuracy and optimizing the runtime efficiency of the K-means initialization. Our findings, for the GTSRB dataset, reveal that starting with 32 or 64 codewords offers comparable accuracy and uncertainty metrics to larger codebooks, while significantly enhancing the efficiency of the K-means initialization. This efficiency highlights the efficacy of our proposed approach.
In addition, for more complex datasets, requiring a broader representation of image features but with minimal initialization time, we employ codeword segmentation to enhance selection capacity efficiently. We explore the impact of segmentation factors of 1, 2, and 4, starting with 16 codewords for GTSRB and 32 for CIFAR100. Our findings indicate that, particularly for CIFAR100, splitting vectors into 4 segments with only 32 initialized codewords achieves impressive performance. Similarly, for GTSRB, segmentation into 2 parts is adequate for effective image feature representation.
[…]
8 $\mathcal{D}_{1}$ 3.61 4.78 2.02 0.257 $8\rightarrow 16\rightarrow 32\rightarrow 64$
$\mathcal{D}_{2}$ 3.68 4.86 2.01 0.265
$\mathcal{D}_{3}$ 3.38 4.81 2.03 0.249
16 $\mathcal{D}_{1}$ 3.41 10.46 1.36 0.515 $16\rightarrow 32\rightarrow 64$
$\mathcal{D}_{2}$ 3.55 10.54 1.37 0.506
$\mathcal{D}_{3}$ 3.23 10.61 1.39 0.486
32 $\mathcal{D}_{1}$ 0.127 25.91 0.0412 0.956 $32\rightarrow 64\rightarrow 128$
$\mathcal{D}_{2}$ 0.0885 25.42 0.0313 0.966
$\mathcal{D}_{3}$ 0.178 26.37 0.117 0.911
64 $\mathcal{D}_{1}$ 0.0975 26.79 0.0086 0.965 $64\rightarrow 128$
$\mathcal{D}_{2}$ 0.0853 26.32 0.0084 0.974
$\mathcal{D}_{3}$ 0.1907 27.23 0.0166 0.926
128 $\mathcal{D}_{1}$ 0.0512 38.73 - 0.954 $128\rightarrow 256$
$\mathcal{D}_{2}$ 0.0453 34.16 - 0.968
$\mathcal{D}_{3}$ 0.0726 43.42 - 0.917
256 $\mathcal{D}_{1}$ 0.0543 41.20 0.0043 0.962 $256\rightarrow 512$
$\mathcal{D}_{2}$ 0.0301 38.96 0.0054 0.959
$\mathcal{D}_{3}$ 0.0577 50.57 0.0103 0.904
Codebook Data $\mathbf{\mathcal{L}_{code}}$ mP mE mA
w/o init $\mathcal{D}_{1}$ 0.6202 6.26 0.125 0.888
$\mathcal{D}_{2}$ 0.6063 5.99 0.296 0.554
$\mathcal{D}_{3}$ 0.6096 5.35 0.267 0.622
w/ init $\mathcal{D}_{1}$ 0.0862 59.64 0.0935 0.945
$\mathcal{D}_{2}$ 0.0785 57.58 0.1604 0.906
$\mathcal{D}_{3}$ 0.0779 99.38 0.1509 0.929
Extensible Codebook v.s. Static Large Codebook. To validate our extensible codebook’s superiority over starting with a large codebook, we ensured both methods ended with the same number of codewords through experiments. For the CIFAR100 dataset, the extensible codebook was initially set to 128 codewords and expanded twice, while the static codebook was fixed at 512 codewords. Results showcased in Table 16 demonstrate the difficulties associated with a larger initial codebook in codeword selection for image features. Conversely, gradually expanding the codebook significantly improved codeword differentiation, yielding better outcomes, such as enhanced accuracy (0.375 for Domain 1) and reduced uncertainty (0.78 vs. 1.66 for the static approach). In addition, perplexity results reveal increased utilization of our extensible codebook, offering clear evidence of our design’s superiority.